# Complexity of deep computations via topology of function spaces

### v3

@@ 1.2. NIP as a universal dividing line between polynomial and exponential complexity
Sauer [SAU72] and Shelah [SHE72] proved the following independently: Let $\mathscr{F}$ be a family of subsets of some set $S$. Either: for every $n\in\mathbb{N}$ there is a set $A\subseteq S$ with $|A|=n$ such that $|\{S_{i}\cap A:i\in\mathbb{N}\}|=2^{n}$ ($\mathscr{F}$ has “exponential complexity”); or: there exists $N\in\mathbb{N}$ such that for every $A\subseteq S$ with $|A|\geq N$, one has
@@
($\mathscr{F}$ has “polynomial complexity”). Clearly, any family $\mathscr{F}$ of subsets of a finite set $S$ has polynomial complexity. The “polynomial” name is justified: indeed, for fixed $N>0$, as a function of the size $|A|=m>0$, we have
@@ 1.3. Rosenthal compacta
Rosenthal compacta possess significant topological and dynamical tameness properties, and play an important role in functional analysis, measure theory, dynamical systems, descriptive set theory, and model theory. In this paper, we use them to study deep computations. For this, we shall first focus on countable languages, which is the theme of the next subsection.
@@
Fix an arbitrary (at most) countable set $\mathcal{P}$ whose elements $P\in\mathcal{P}$ will be called predicate symbols or formal predicates. Our present goal is to characterize relatively compact subsets of $B_{1}(X,\mathbb{R}^{\mathcal{P}})$, where $X$ is always assumed to be a perfectly normal space (often a Polish space).
@@
$\lambda_{x}:\mathbb{R}^{X\times\mathcal{P}}\to\mathbb{R}^{\mathcal{P}}:\varphi\mapsto\lambda_{x}(\varphi)\coloneqq\varphi(x,\cdot);$
@@
$\rho_{P}:\mathbb{R}^{X\times\mathcal{P}}\to\mathbb{R}^{X}:\varphi\mapsto\varphi(\cdot,P).$
@@
$\pi_{x}:\mathbb{R}^{X}\to\mathbb{R}:f\mapsto f(x),$ $\pi_{P}:\mathbb{R}^{\mathcal{P}}\to\mathbb{R}:f\mapsto f(P).$
@@
Lastly, a simple but useful lemma that helps understand when we restrict a set of functions to a specific subspace of the domain space, we may always assume that the subspace is closed, as replacing the subspace by its closure has no effect on NIP.
@@
2. Compositional Computation Structures: A structural approach to arbitrary-precision arithmetic
In this section, we connect function spaces with arbitrary-precision arithmetic computations. We start by summarizing some basic concepts from [ADI+24].
A computation states structure is a pair $(L,\mathcal{P})$, where $L$ is a set whose elements we call states and $\mathcal{P}$ is a collection of real-valued functions on $L$ that we call predicates. For a state $v\in L$, the type of a state $v$ is the indexed family
$\operatorname{tp}(v)=(P(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}.$
For each $P\in\mathcal{P}$, we call the value $P(v)$ the $P$-th feature of $v$. A transition of a computation states structure $(L,\mathcal{P})$ is a map $f:L\to L$.
Intuitively, $L$ is the set of states of a computation, and the predicates $P\in\mathcal{P}$ are primitives that are given and accepted as computable. Each state $v\in L$ is uniquely characterized by its type $\operatorname{tp}(v)$, so we may identify $L$ with a subset of $\mathbb{R}^{\mathcal{P}}$. Important state spaces are $L=\mathbb{R}^{\mathbb{N}}$ and $L=\mathbb{R}^{n}$ for some positive integer $n$, endowed with predicate $P_{i}(v)=v_{i}$, one each for the $i$-th coordinate of $v$. We regard the space of types as a topological space, endowed with the topology of pointwise convergence induced by the product topology of $\mathbb{R}^{\mathcal{P}}$. Via the identification $v\mapsto\operatorname{tp}(v)$, the states space $L$ is correspondingly topologized; in particular, for each $P\in\mathcal{P}$, the projection map $v\mapsto P(v)$ is continuous.
@@ 2. Compositional Computation Structures: A structural approach to arbitrary-precision arithmetic
Given a computation states structure $(L,\mathcal{P})$, any element of $\mathbb{R}^{\mathcal{P}}$ in the image of $L$ under the map $v\mapsto\operatorname{tp}(v)$ will be called a realized (state) type. The topological closure of the set of realized types in $\mathbb{R}^{\mathcal{P}}$ (endowed with the pointwise convergence topology) will be called the space of types of $(L,\mathcal{P})$, denoted $\mathcal{L}$; elements $\xi\in\mathcal{L}$ are called state types. Elements of $\mathcal{L}\setminus L$ will be called unrealized types.
Intuitively, state types capture a notion of “limit state”.
As we combine ideas of model theory [KEI03] and topology [BFT78], we are interested in families of real-valued functions that are pointwise bounded. This leads us to the concepts of sizer and shard introduced first in introduced in [ADI+24]:
@@
A sizer is a family $r_{\bullet}=(r_{P})_{P\in\mathcal{P}}$ of positive real numbers, indexed by $\mathcal{P}$. Given a sizer $r_{\bullet}$, let $\mathbb{R}^{[r_{\bullet}]}=\prod_{P\in\mathcal{P}}[-r_{P},r_{P}]$ (a compact space), and let the $r_{\bullet}$-shard of a states space $L$ be
$L[r_{\bullet}]=L\cap\mathbb{R}^{[r_{\bullet}]}.$
For a sizer $r_{\bullet}$, the $r_{\bullet}$-type shard is defined as $\mathcal{L}[r_{\bullet}]=\overline{L[r_{\bullet}]}$ (a closed, hence compact subset of $\mathbb{R}^{[r_{\bullet}]}$).
Let also $\mathcal{L}_{\text{sh}}$ be the union of all type-shards as the sizer $r_{\bullet}$ varies.
In general, $\mathcal{L}_{\text{sh}}\subseteq\mathcal{L}$, and the inclusion may be proper. However, equality holds in the important special case when $\mathcal{P}$ is countable (see [ADI+24]).
@@ A Compositional Computation Structure (CCS) is a triple $(L,\mathcal{P},\Gamma)$, where
Elements of the semigroup $\Gamma$ are called the computations of the structure $(L,\mathcal{P},\Gamma)$. We assume that the identity map id on $L$ is an element of $\Gamma$ (which is thus not merely a semigroup but a monoid of transformations of $L$).
We topologize $\Gamma$ as a subset of the topological product $L^{L}$, where the “exponent” $L$ serves merely as an index set, but the “base” $L$ is topologized by type; consequently, one may identify $\Gamma$ with a subset of the topological product $(\mathbb{R}^{\mathcal{P}})^{L}$. More specifically, $\Gamma$ is identified with a subset of $\mathcal{L}^{L}$, which is a closed subspace of $(\mathbb{R}^{\mathcal{P}})^{L}$. Therefore, we have an inclusion $\overline{\Gamma}\subseteq\mathcal{L}^{L}$. Elements $\xi\in\overline{\Gamma}$ are called (real-valued) deep computations or ultracomputations.
@@
2.2. Computability and definability of deep computations and the Extendibility Axiom
Let $f:L\to\mathcal{L}$ be a function that maps each input state type $(P(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$ to an output state type $(P\circ f(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$.
@@ 2.2. Computability and definability of deep computations and the Extendibility Axiom
    We will say that $f$ is definable if for each $Q\in\mathcal{P}$, the output feature $Q{\circ}f:L\to\mathbb{R}$ is a definable predicate in the following sense: There is an approximating function $\varphi_{Q,\,K,\varepsilon}:L\to\mathcal{R}$ that can be built recursively out of a finite number of the (primitively computable) predicates in $\mathcal{P}$ and by a finite number of iterations of the finitary lattice operations $\land$ (=min) and $\lor$ (=max), the operations of $\mathbb{R}^{\mathbb{R}}$ as a vector algebra (that is, vector addition and multiplication and scalar multiplication) and the operators sup and inf applied on individual variables from $L$, and such that
    $|\varphi_{Q,\,K,\varepsilon}(v)-Q{\circ}f(v)|\leq\epsilon,\qquad\text{for all $v\in K$}.$
    Remark: What we have defined above is a model-theoretic concept; it is a special case of the concept of first-order definability for real-valued predicates in the model theory of real-valued structures first introduced in [IOV94] for model theory of functional analysis and now standard in model theory (see [KEI03]). The $\land$ (=min) and $\lor$ (=max) operations correspond to the positive Boolean logical connectives “and” and “or”, and the sup and inf operators correspond to the first-order quantifiers, $\forall$ and $\exists$.
@@
    We will say that $f$ is computable if it is definable in the sense defined above under (1), but without the use of the sup/inf operators; in other words, if for every choice of $Q,K,\varepsilon$, the approximation function $\varphi_{Q,K,\varepsilon}:L\to\mathcal{L}$ can be constructed without any use of sup or inf operators. This is quantifier-free definability (i.e., definability as given by the preceding paragraph, but without use of quantifiers), which, from a logic viewpoint, corresponds to computability (the presence of the quantifiers $\exists$ and $\forall$ are the reason behind the undecidability of first-order logic).
It is shown in [ADI+24] that:
@@
    For a definable $f:L\to\mathcal{L}$, the approximating functions $\varphi_{Q,\,K,\varepsilon}$ may be taken to be polynomials of the input features, and
@@
    Definable transforms $f:L\to\mathcal{L}$ are precisely those that extend to continuous $\tilde{f}$$\mathcal{L}\to\mathcal{L}$.
To summarize, a function $f:L\to\mathcal{L}$ is computable if and only if it is definable if and only if it is polynomially approximable if and only it can be extended to a continuous $\tilde{f}$$\mathcal{L}\to\mathcal{L}$. This is the reason for the following definition.
Definition 2.5.
We say that a CCS $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom if for all $\gamma\in\Gamma$, there is $\tilde{\gamma}:\mathcal{L}_{\text{sh}}\to\mathcal{L}_{\text{sh}}$ such that for every sizer $r_{\bullet}$ there is a sizer $s_{\bullet}$ such that $\tilde{\gamma}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\to\mathcal{L}[s_{\bullet}]$ is continuous. We refer to $\tilde{\gamma}$ as a free extension of $\gamma$.
For the rest of the paper, fix for each $\gamma\in\Gamma$ a free extension $\tilde{\gamma}$ of $\gamma$. For any $\Delta\subseteq\Gamma$, let $\tilde{\Delta}$ denote $\{\tilde{\gamma}\colon\gamma\in\Delta\}$.

### v4

@@ 1.2. NIP as a universal dividing line between polynomial and exponential complexity
Sauer [SAU72] and Shelah [SHE72] proved the following independently: Let $\mathscr{F}$ be a collection of subsets of some set $S$. Either: for every $n\in\mathbb{N}$ there is a set $A\subseteq S$ with $|A|=n$ such that $|\{S_{i}\cap A:i\in\mathbb{N}\}|=2^{n}$ ($\mathscr{F}$ is a collection having “exponential complexity”); or: there exists $N\in\mathbb{N}$ such that for every $A\subseteq S$ with $|A|\geq N$, one has
@@
($\mathscr{F}$ is a collection having “polynomial complexity”). Clearly, any collection $\mathscr{F}$ of subsets of a finite set $S$ has polynomial complexity. The “polynomial” name is justified: indeed, for fixed $N>0$, as a function of the size $|A|=m>0$, we have
@@ 1.3. Rosenthal compacta
Rosenthal compacta possess significant topological and dynamical tameness properties, and play an important role in functional analysis, measure theory, dynamical systems, descriptive set theory, and model theory. In this paper, we use them to study deep computations.
@@
Fix an arbitrary (at most) countable set $\mathcal{P}$ whose elements $P\in\mathcal{P}$ will be called predicate symbols or formal predicates. Our present goal is to characterize relatively compact subsets of $B_{1}(X,\mathbb{R}^{\mathcal{P}})$, where $X$ is always assumed to be a perfectly normal space (typically a Polish space).
@@
$\lambda_{x}:\mathbb{R}^{X\times\mathcal{P}}\to\mathbb{R}^{\mathcal{P}}:\varphi\mapsto\lambda_{x}(\varphi)\coloneqq\hat{}\varphi(x);$
@@
$\rho_{P}:\mathbb{R}^{X\times\mathcal{P}}\to\mathbb{R}^{X}:\varphi\mapsto\rho_{P}(\varphi)\coloneqq\varphi\,\hat{}(P).$
@@
$\lambda_{x}:\mathbb{R}^{X}\to\mathbb{R}:f\mapsto f(x),$ $\rho_{P}:\mathbb{R}^{\mathcal{P}}\to\mathbb{R}:f\mapsto f(P).$
@@
Lastly, a simple but useful lemma to the effect that a set of continuous functions has NIP over a subset of the domain if and only if it has NIP over the closure of the subset.
@@
Historical remarks
The Baire hierarchy of functions was introduced by René-Louis Baire in his 1899 doctoral thesis, Sur les fonctions de variables réelles. His work moved away from the 19th-century preoccupation with “pathological” functions toward a constructive classification based on pointwise limits.
$C_{p}$-theory is an active field of research in general topology pioneered by A. V. Arhangel’skiĭ and his students in the 1970s and 1980s [ARK92]. $C_{p}$-theory has found many applications in model theory and functional analysis. For a recent survey, see [TKA11].
2. Compositional Computation Structures:
A structural approach to arbitrary-precision computations
In this section, we connect function spaces with arbitrary-precision computations. We start by summarizing some basic concepts from [ADI+24].
2.1. Computation States Structures
A computation states structure is a pair $(L,\mathcal{P})$, where $L$ is a set whose elements we call states and $\mathcal{P}$ is a collection of real-valued functions on $L$ that we call predicates, also called features. For each $P\in\mathcal{P}$, we call the value $P(v)$ the $P$-th (observable) feature of $v$.
A transition of a computation states structure $(L,\mathcal{P})$ is a map $f:L\to L$.
The type of a state $v\in L$ is the indexed family [^5] As in §1.10 above, we regard each predicate on $L$ as named by a specific formal label $P$ (a predicate symbol) which is given the semantic meaning of a bona fide real function $P(\cdot):v\mapsto P(v)$ (the predicate proper) in the computation states structure. Such formal symbol $P$ becomes de facto the name of an observable feature having the real value $P(v)$ at any state $v\in L$. Whenever $\mathcal{P}$ is used as an index set (e.g., as in the topological product $\mathbb{R}^{\mathcal{P}}$ below), it is regarded as a set of formal symbols. This syntactic vs. semantic abuse of nomenclature is commonplace in mathematics, but the model-theoretic framework requires contextual awareness of the distinction. To disambiguate when necessary, we shall write $P(\cdot)$, implying the semantic meaning of a real function on $L$, in contrast to $P$, meant as a purely syntactic symbol. Nevertheless, by a standard abuse of nomenclature in appropriate contexts, we may also treat $P$ as a function on states (i.e., take $P$ to mean $P(\cdot)$ proper).
$\operatorname{tp}(v)=(P(v))_{P\in\mathcal{P}}\in\mathbb{R}^{\mathcal{P}}$
of its features.
Intuitively, $L$ is the set of states of a computation, and the predicates $P\in\mathcal{P}$ are primitives that are given and accepted as computable. (I.e., each symbol $P\in\mathcal{P}$ denotes a function $P(\cdot)$ taken as a computational primitive on $L$.)
We assume that each state $v\in L$ is uniquely characterized by its type $\operatorname{tp}(v)$, so we may identify $L$ with a topological subspace of $\mathbb{R}^{\mathcal{P}}$ and topologized as such (i.e., $L$ is identified with its forward image $\operatorname{tp}[L]\subseteq\mathbb{R}^{\mathcal{P}}$ under the type map). As a topological space, $L$ is thus endowed with the topology of pointwise convergence induced by the product topology of $\mathbb{R}^{\mathcal{P}}$. In particular, for each $P\in\mathcal{P}$, the projection map $v\mapsto P(v)$ is continuous.
Important state spaces include $L=\mathbb{R}^{\mathbb{N}}$ (resp., $L=\mathbb{R}^{n}$ for some positive integer $n$), endowed with one predicate $P_{i}(v)=v_{i}$ (giving the $i$-th coordinate of $v$) for each $i\in\mathbb{N}$ (resp., for each $0\leq i<n$). State subspaces $L\subseteq\mathbb{R}^{\mathbb{N}}$ and $L\subseteq\mathbb{R}^{n}$ are of primary interest in this paper. (Any states structure having a finite or countable family of predicates is effectively such a subspace.)
@@ 2.1. Computation States Structures
Given a computation states structure $(L,\mathcal{P})$, any type $\operatorname{tp}(v)\in\mathbb{R}^{\mathcal{P}}$ of a state $v\in L$ will be called a realized (state) type. The topological closure of the set of realized types in $\mathbb{R}^{\mathcal{P}}$ (endowed with the pointwise convergence topology) will be called the space of (state) types of $(L,\mathcal{P})$, denoted $\mathcal{L}$; elements $\mathbf{v}\in\mathcal{L}$ are called state types. Elements $\mathbf{v}\in\mathcal{L}\setminus L$ are unrealized state types.
Intuitively, state types capture a notion of “limit state”. For each $P\in\mathcal{P}$, the $P$-th coordinate map (natural projection) $\mathbb{R}^{\mathcal{P}}\to\mathbb{R}:\mathbf{v}=(\mathbf{v}_{Q})_{Q\in\mathcal{P}}\mapsto\mathbf{v}_{P}$, gives a continuous extension to $\mathcal{L}$ of the predicate $P(\cdot)$ on $L$; abusing notation, the extension will still be denoted $P(\cdot)$ (or just $P$).
Remark 2.2.
Just as $L$ is identified with a subset of $\mathcal{L}$, but going in the opposite direction, one may regard unrealized state types as “hyper-states” (when such type is obtained as a sequential limit—or ultralimit—of realized state types, unrealized state types may be called “deep states”). A comment for readers versed in model theory: at least for the purpose of realizing state types (with parameters in the original states space $L$), the type space $(\mathcal{L},\mathcal{P})$ serves as a “poor man’s” ultraproduct of $(L,\mathcal{P})$ wherein, by the very definition of $\mathcal{L}$, each type $\mathbf{v}$ is realized.
2.1.1. Bounds on precision and magnitude. Shards
In order for computation state structures to serve their purpose as models of real-world computing, we must keep in mind that computations involving real-valued quantities can only ever carried out approximately, to a preset degree of precision (i.e., to within acceptable error bounds). More subtly, yet equally important in practice, one must also bound the magnitude of real quantities involved in all computations: one cannot possibly store—let alone compute with—arbitrarily large numbers! Thus, although the collection $\mathcal{P}$ of predicates may be infinite (possibly, uncountable even), “true” computations must imply choices, at the outset, of bounds for errors and for absolute magnitudes of quantities involved. Heuristically, different parts of a computation may require suitably matching (different) error bounds, as well as dealing with quantities of very different sizes. Focusing for now on the second of these issues, our results shall generally apply to sets of computations for which arbitrary but fixed bounds for each predicate are set in advance. This leads to the concept of shard first introduced in [ADI+24], as per the following definition.
@@
Fix a computation states structure $\underline{L}=(L,\mathcal{P})$, with $L\subseteq\mathbb{R}^{\mathcal{P}}$ implicitly identified with its type space. A sizer is a family $r_{\bullet}=(r_{P})_{P\in\mathcal{P}}$ of positive real numbers, indexed by $\mathcal{P}$. Given a sizer $r_{\bullet}$, let $\mathbb{R}[r_{\bullet}]\coloneqq\prod_{P\in\mathcal{P}}[-r_{P},r_{P}]$ (a compact space), and let the $r_{\bullet}$-shard of the states space $L$ be
$L[r_{\bullet}]=L\cap\mathbb{R}[r_{\bullet}].$
For a sizer $r_{\bullet}$, the $r_{\bullet}$-type shard is defined as $\mathcal{L}[r_{\bullet}]=\overline{L[r_{\bullet}]}$ (a closed, hence compact subset of $\mathbb{R}[r_{\bullet}]$).
Let also $\mathcal{L}_{\text{sh}}$ (the space of shard-supported types) be the union of all type-shards as the sizer $r_{\bullet}$ varies.
Evidently, $\mathcal{L}_{\text{sh}}\subseteq\mathcal{L}$, although the inclusion may be proper. However, the equality $\mathcal{L}_{\text{sh}}=\mathcal{L}$ holds in the important special case when $\mathcal{P}$ is (at most) countable (see [ADI+24]).
@@ A Compositional Computation Structure (CCS) is a triple $(L,\mathcal{P},\Gamma)$, where
Elements of the semigroup $\Gamma$ are called the computations of the structure $(L,\mathcal{P},\Gamma)$. The state space $L$ is implicitly identified with a subset of the type space $\mathcal{L}\subseteq\mathbb{R}^{\mathcal{P}}$ of $(L,\mathcal{P})$. We also assume that the identity map id on $L$ is an element of $\Gamma$ (which is thus not merely a semigroup but a monoid of transformations of $L$).
We topologize $\Gamma$ as a subset of the topological product $L^{L}$, where (as usual) the “exponent” $L$ serves merely as an index set, but the “base” $L\subseteq\mathbb{R}^{\mathcal{P}}$ has the relative product topology (of pointwise convergence). In this manner, the semigroup $\Gamma$ is identified with a subset of the topological product $\mathcal{L}^{L}\subseteq(\mathbb{R}^{\mathcal{P}})^{L}$, which leads to an inclusion $\overline{\Gamma}\subseteq\mathcal{L}^{L}$. Elements $\xi\in\overline{\Gamma}$ are called (real-valued) deep computations or ultracomputations.
@@
Unlike its subspace $L^{L}$ (a topological semigroup), the space $\mathcal{L}^{L}$ does not have a natural semigroup structure, so ultracomputations are not generally composable. Intuitively, ultracomputations are “final” or “ultimate”.
@@
Remarks 2.6.
(For readers versed in model theory.) [^6] Refer to [KEI23] where a general framework for real-valued structures is introduced; see also [DI17, section 6] for a basic tutorial.
@@
    Compositional computational structures as defined above are not real-valued structures in a literal sense, because there is no intrinsic structural manner to ensure that $\Gamma$ is realized as a literal subset of the product space $L^{L}$. However, in a semantically transparent manner, any CSS as above may alternatively be seen (per its original definition in [ADI+24]) as a real-valued structure $(L,\mathcal{P},\Gamma,\circ,\operatorname{ev})$, where
    • $(L,\mathcal{P})$ is a computation states structure,
    • $(\Gamma,\circ)$ is a semigroup,
    • $\operatorname{ev}:\Gamma\times L\to L:(\gamma,v)\mapsto\operatorname{ev}(\gamma,v)$ is a semigroup action of $\Gamma$ on $L$ (i.e., $\operatorname{ev}(\gamma\circ\delta,v)=\operatorname{ev}(\gamma,\operatorname{ev}(\delta,v))$ for $\gamma,\delta\in\Gamma$ and $v\in L$).
    Both definitions are reconciled upon identifying $\gamma\in\Gamma$ with the map $(\gamma,v)\mapsto\operatorname{ev}(\gamma,v)$.
@@
    In the preceding sense of real-valued structures, the class of compositional computation structures (for a fixed predicate symbol collection $\mathcal{P}$) is first-order elementary. The limited notion of state type we have introduced (essentially, quantifier-free state types with parameters of sort $L$ only) may be enriched to state types involving more formulas (e.g., with parameters in $\Gamma$, or including quantified formulas), and extended to “transform-types” (i.e., types of sort $\Gamma$). In particular, given a structure $\underline{C}=(L,\mathcal{P},\Gamma,\circ,\operatorname{ev})$, all types are realized in some ultrapower $\underline{C}^{(\mathcal{U})}$ of $\underline{C}$ (constructed in the usual manner detailed in the references above). [^7] By contrast (cf., Remark 2.2), the ultrapower construction is not necessary to realize state types in a computation states structures. This approach to spaces of types was introduced in Banach space theory by Krivine [KRI74, KM81, KRI72].
2.3. Computability of deep computations and the Extendibility Axiom
Let $(L,\mathcal{P})$ be a computation structure with countable predicate collection $\mathcal{P}$.
@@ 2.3. Computability of deep computations and the Extendibility Axiom
    A computable predicate is any real function $\varphi:L\to\mathbb{R}$ whose restriction to an arbitrary shard is uniformly approximable by polynomials (with real coefficients) in the features $P(\cdot)$.
@@
    A transform $f:L\to\mathcal{L}$ is computable if, for each $Q\in\mathcal{P}$, the output feature $Q{\circ}f:L\to\mathbb{R}$ is a computable predicate.
It is shown in [ADI+24] that computable transforms $f:L\to\mathcal{L}$ are precisely those that extend to continuous maps $\tilde{f}$$\mathcal{L}\to\mathcal{L}$.
To summarize, for a function $f:L\to\mathcal{L}$, the following conditions are equivalent:
  • $f$ is computable
  • $f$ has is uniformly polynomially approximable features on shards
  • $f$ extends to a continuous map $\mathcal{L}\to\mathcal{L}$.
These equivalences motivate the following definition:
A computation $\gamma\in\Gamma$ is extendable if it has some extension $\tilde{\gamma}:\mathcal{L}_{\text{sh}}\to\mathcal{L}_{\text{sh}}$ such that, for every sizer $r_{\bullet}$, there is a sizer $s_{\bullet}$ such that $\tilde{\gamma}|_{\mathcal{L}[r_{\bullet}]}:\mathcal{L}[r_{\bullet}]\to\mathcal{L}[s_{\bullet}]$ is continuous.
Because $\tilde{\gamma}$ extends $\gamma$ continuously on each compact type-shard, we call it the $K$-extension of $\gamma$; such extension is evidently unique (if it exists).
Abusing notation, for any $\Delta\subseteq\Gamma$ consisting of extendable computations, the set $\{\tilde{\gamma}\colon\gamma\in\Delta\}$ will be denoted $\tilde{\Delta}$.
Extendibility Axiom.
Every computation $\gamma\in\Gamma$ is extendable.
@@
For the rest of the paper, the Extendibility Axiom is imposed on all CSSs.